a tortoise can run with a speed of 0.10 m/s and a hare can run 20 times as fast. in a race, they both start at the same time, but the hare stops to rest for 2minutes. The tortoise wins by a shell (20cm). (a) how long does the race take? (b) what is the length of the race?.
1
Expert's answer
2017-07-28T15:54:07-0400
(a)
d=2t+0.2.
d=0.1(t+120).
2t+0.2=0.1(t+120)
t=6.21 s.
(b)
d=2(6.21)+0.2=12.62 m.